ROTATIONALLY RESOLVED Ã^{2}A_{1} - X̃^{2}E ELECTRONIC SPECTRA OF SYMMETRIC METHOXY RAD-ICALS: CH_{3}O AND CD_{3}O

\underline{MING-WEI CHEN}, JINJUN LIU,^{a} DMITRY G. MELNIK, and TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus, Ohio 43210.

Methoxy radical, a key component in both atmospheric and combustion chemistry, has attracted spectroscopic interest for more than twenty years. Microwave measurements of CH_{3}O^{b} and CD_{3}O^{c} with precision on the kHz scale have determined the X̃^{2}E parameters. Jet-cooled laser induced fluorescence (LIF) spectra^{d} have also been observed by our group for both CH_{3}O and CD_{3}O, with high-resolution (Δν ≈250 MHz) and high-accuracy (Δσ ≈50 MHz), for the 3_{0}^{2} and 6_{0}^{1} bands of the Ã^{2}A_{1} - X̃^{2}E_{3/2} electronic transition. Since the ground state component E_{1/2} is ≈60 cm^{−1} energetically higher than the E_{3/2} spin component, the X̃^{2}E_{1/2} state is not thermally populated in a jet-cooled environment. However, our complementary stimulated emission pumping (SEP) experiment directly interrogates the X̃^{2}E_{1/2} level of CH_{3}O and CD_{3}O by depleting the fluorescence from Ã^{2}A_{1} 3^{2} excited levels. It has now been performed with the same resolution and accuracy as the LIF work. The global analysis of the microwave^{b,c}, LIF, and SEP data breaks correlations in the microwave data and provides better determinations for the X̃ and Ã states’ parameters. Comparison of the values for CH_{3}O, ^{13}CH_{3}O^{e}, and CD_{3}O allows us to separate first-order from second-order electronic and vibrational contributions based upon the isotopic dependencies of the effective ground state parameters, e.g. the Jahn-Teller parameters h_{1} and h_{2}, spin-rotation parameters, etc.

apresent address: Laboratory of Physical Chemistry, ETH Zurich, Wolfgang-Pauli-Str. 10, 8093 Zurich, Switzerland.

bY. Endo, S. Saito, and E. Hirota, J. Chem. Phys., \underline{81}, 122 (1984)

cT. Momose, Y. Endo, and E. Hirota, unpublished data.

dJ. Liu, J. T. Yi, V. Stakhursky, and T. A. Miller TJ04 and 05 61^{st} International Symposium on Molecular Spectroscopy, (2006)

eT. Momose, Y. Endo, E. Hirota, and T. Shida, J. Chem. Phys., \underline{88}, 5338 (1988)